Univalence of certain integral transforms S. Ponnusamy, Helsinki, Finland, and Vikramaditya Singh, Azad Nagar, India

Keywords: Univalent, starlike, convex and close-to-convex functions AMS Subject classification: 30C45 Abstract: This paper finds sufficient conditions for univalence of certain class of functions and of certain integral operators, respectively.